OpenStudy (anonymous):

How do I turn 3x -3y = 15 into equivalent form?

OpenStudy (tkhunny):

Have you considered algebraic operations? Divide by 15 to produce the Intercept Form \(\dfrac{x}{5}+\dfrac{y}{-5} = 1\)

OpenStudy (anonymous):

Im sorry, how did you get that?

OpenStudy (tkhunny):

Already answered. Divide by 15. Subtract 15 to obtain Normal Form 3x -3y - 15 = 0